OpenStudy (vshiroky):

Find the equation of the line that contains the given point and is perpendicular to the given point and is perpendicular to the given line. Write the equation in slope-intercept form, if possible. (1,-1); y=-x+3 What is the equation of the line?

OpenStudy (***[isuru]***):

when 2 lines r perpendicular to each other if the slope of one lie is "m" then the slope of the other is "(-1/m)" so... in y= -x + 3... the slope is -1 which means a line perpendicular to that will have the slope (+1) so... u can write the equation as y= x + b where b = intercept now to find intercept u can use ( 1 , -1) if u plug them in to the equation -1 = 1 + b b = -2 so .. the equation of the line is y= x - 2 hope this will help ya:)